how many quarts of type P punch, which is 3% alcohol and type Q punch, which is 8% alcohol, should be combined to get 20 quarts of punch that is 6% alcohol?
Found 2 solutions by stanbon, Boreal:
Answer by stanbon(75887)

how many quarts of type P punch, which is 3% alcohol and type Q punch, which is 8% alcohol, should be combined to get 20 quarts of punch that is 6% alcohol?
------------
Equation:
alch + alch = alch
----
0.03p + 0.08(20-p) = 0.06*20
-------
3p + 8*20 - 8p = 6*20
----
-5p = -2*20
-----
p = 8 qts (amt. of P punch needed)
20-p = 12 qts (amt of Q punch needed)
============
Cheers,
Stan H.

Answer by Boreal(15235)

p=P punch in quarts
q=Q punch in quarts
Because the final amount is closer to Q than to P, expect Q to be larger.
.03P+ .08(20-P)=.06 *20 =1.2 alcohol-quarts unit
.03P +1.6-.08P=1.2
-.05P=-.4
divide both sides by -.05
P=8 quarts
Q=12 quarts
.03*8 + .08*12= .24+.96=1.2